OpenStudy (anonymous):

(MEDALS!!!!!!) Please answer. Find the probabilities. Write your answer as a fraction, as a decimal, and as a percent. If necessary, round decimals to the nearest thousandth and percents to the nearest tenth. 1. A bag contains 15 red beads, 15 white beads, 20 green beads, and 25 blue beads. What is the probability of randomly drawing a blue bead?

OpenStudy (anonymous):

The total probability space is the sum of all 75 The probability of getting a blue is the number of blue beads divided by the total probability space 25/75 = 1/3 = 0.333333 = 33.33333%